# Exploring Cross-model Neuronal Correlations in the Context of Predicting Model Performance and Generalizability

### v5

@@ Methods
Additionally, we explore the vulnerability of neural networks to black-box adversarial attacks and examine how this susceptibility relates to our proposed metric. To investigate the architectural relationships between different neural networks, we adapted the attack method introduced by (Papernot et al. 2017), with slight modifications to their original approach. This black-box method facilitates model comparison without necessitating prior knowledge of their architectures or training datasets. As described in Algorithm 1, at each iteration, we train a simple model like a two-layer CNN that differs from the other models used in the experiments on the dataset. Subsequently, we adjust the data using the perturbation computed via the Fast Gradient Sign Method (FGSM) (Liu, Yu, and Su 2019). We iterate through the process multiple times to generate our ultimate adversaries. In our method, we solely utilize test data instead of custom-crafted data, as the former remains untouched during training. Furthermore, we refrain from altering labels based on the oracle’s feedback and discard intermediate adversaries generated throughout the process, retaining only the final iteration to reduce the time and space complexity.
@@
[…]
@@
(a) MNIST
Net1 Net2 Net3 Net4 Net5
Net1 1 0.644 0.646 0.642 0.576
Net2 X 1 0.698 0.699 0.581
Net3 X X 1 0.733 0.575
Net4 X X X 1 0.572
Net5 X X X X 1
(b) Fashion-MNIST
Net1 Net2 Net3 Net4 Net5
Net1 1 0.200 0.207 0.214 0.167
Net2 X 1 0.278 0.301 0.208
Net3 X X 1 0.289 0.204
Net4 X X X 1 0.221
Net5 X X X X 1
(c) Kuzushiji-MNIST
Net1 Net2 Net3 Net4 Net5
Net1 1 0.564 0.577 0.590 0.523
Net2 X 1 0.613 0.620 0.527
Net3 X X 1 0.671 0.544
Net4 X X X 1 0.510
Net5 X X X X 1
(d) Afro-MNIST
Net1 Net2 Net3 Net4 Net5
Net1 1 0.621 0.613 0.628 0.465
Net2 X 1 0.700 0.712 0.551
Net3 X X 1 0.736 0.556
Net4 X X X 1 0.547
Net5 X X X X 1
@@
A. Small neural networks
We evaluated our method on five simple fully connected networks (FCNs) and a two-layer convolutional neural network (CNN). The number of neurons considered for each hidden layer is described in Table 1. These networks were trained on four datasets including MNIST (LeCun 1998), Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017), Kuzushiji-MNIST (Clanuwat et al. 2018), and Afro-MNIST (Wu, Yang, and Prabhu 2020). All these datasets are comprised of 60000 training examples and 10000 test examples. Each example consists of a $28\times 28$ image paired with a classification label from a set of 10 categories. MNIST, Fashion-MNIST, Kuzushiji-MNIST, and Afro-MNIST contain images of handwritten numerals, fashion, Hiragana characters, and African numbers respectively. The selection of datasets with similar attributes was deliberate as our focus was on understanding the interplay between networks, irrespective of data size or category.
@@ A. Small neural networks
Table 4 depicts the robustness of FCNs against the black-box attack detailed in Algorithm 1. The impact of the attack on some test images is illustrated in Figure 1. Results for CNN are also provided, showcasing its inferior robustness, given its role in generating the adversaries. In our specific scenario, we selected a norm value of $\epsilon=\frac{10}{255}$ and $N_{ep}=10$. The results indicate that, on average, Net5, Net2, Net3, Net4, and Net1 ranked from vulnerable to robust. Among the FCNs, Net5 exhibited the lowest overall accuracy. Notably, according to Table 2(d), Net5 displayed the weakest correlation with the other networks, aligning with its subpar performance and larger number of layers. Conversely, Net4 demonstrated the highest average correlation score with the other networks, potentially attributed to its larger number of neurons per layer.
@@
B. Large neural networks
The proposed correlation metric was also investigated on large neural networks such as ResNet (He et al. 2016) and DenseNet (Huang et al. 2017). ResNet, short for Residual Network, is a deep neural network architecture that leverages residual blocks and skip connections to enhance training efficiency and performance. The variants of ResNet include ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 (He et al. 2016). DenseNet, short for Dense Convolutional Network, is a deep neural network architecture that leverages dense connections between layers to promote feature reuse and efficient gradient flow, resulting in compact models that excel in various computer vision tasks. The architecture has several variants, namely DenseNet-121, DenseNet-161, DenseNet-169, and DenseNet-201 (Huang et al. 2017). For both cases, we utilized the pretrained weights on ImageNet (Deng et al. 2009) as our baselines. We limited our analysis to just 10 test data points sampled from the validation set of ImageNet due to the time and space complexities inherent in our methodology. Since investigating the correlation on all of the neurons in all layers was not possible, we focused our analysis solely on the initial output of the fourth layer for all ResNets and the primary output of the third transition layer for all DenseNets. A point to be highlighted is that assessing the correlation metric on the final layers often yields superior accuracy compared to other layers, given that these layers encapsulate more profound and meaningful representations. The results shown in Table 5 suggest that the most similar networks to ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 are ResNet-34, ResNet-18, ResNet-34, ResNet-34, and ResNet-34 respectively. As demonstrated in Table 6, DenseNet-161, DenseNet-121, DenseNet-201, and DenseNet-121 have the most similarities to DenseNet-121, DenseNet-161, DenseNet-169, and DenseNet-201. This pattern illustrates that networks with roughly the same number of layers often exhibit more profound partial correlations, serving as a validation of the effectiveness of our proposed correlation metric. Furthermore, deeper neural networks exhibit greater robustness against adversarial attacks. Thus, our findings further reveal that a stronger correlation with a robust model correlates with higher robustness.
@@ B. Large neural networks
Discussion
@@
Conclusion
In the ongoing effort to establish a robust framework for AI trustworthiness, a critical component of any eventual solution will be analyses that can be performed by objective and independent researchers without the need for the data the model was trained with initially. One possible approach may be to examine the architecture of the model in question in reference to another established model, in the hopes of gleaning insights into robustness and generalizability. We proposed an approach of taking an aggregated correlation score of neuronal activation across two independent neural networks. Our resulting correlation score effectively reveals connections across diverse neural networks. This score highlights performance implications, suggesting that lower correlations may signal deviations from expected activation patterns, and further indicates that strong correlations with robust models imply similar expectations for performance in other networks. Additionally, our method addresses a critical gap in responsible AI development by providing a robust approach to evaluate model quality and generalizability, independent of internal developer datasets. In spite of these promising results, this approach faces limitations in precision with larger models and cannot pinpoint particular reasons for low correlation scores, making inferences less specific. Future research should focus on resolving the deficiencies, optimizing efficiency, and enhancing diagnostic capabilities to deepen insights into neural network behaviors, ultimately supporting the safe and effective deployment of AI systems.

### v7

@@ 3.5 Black-Box Adversarial Robustness
Additionally, we explore the vulnerability of neural networks to black-box adversarial attacks and examine how this susceptibility relates to our proposed metric. To investigate the architectural relationships between different neural networks, we adapted the attack method introduced by (blackbox), with slight modifications to their original approach. This black-box method facilitates model comparison without necessitating prior knowledge of their architectures or training datasets. As described in Algorithm 1, at each iteration, we train a simple model like a two-layer CNN that differs from the other models used in the experiments on the dataset. Subsequently, we adjust the data using the perturbation computed via the Fast Gradient Sign Method (FGSM) fgsm. We iterate through the process multiple times to generate our ultimate adversaries. In our method, we solely utilize test data instead of custom-crafted data, as the former remains untouched during training. Furthermore, we refrain from altering labels based on the oracle’s feedback and discard intermediate adversaries generated throughout the process, retaining only the final iteration to reduce the time and space complexity.
@@ 4 Empirical Results
        Refer to caption Refer to caption Refer to caption Refer to caption                     Refer to caption Refer to caption Refer to caption Refer to caption                  Refer to caption Refer to caption Refer to caption Refer to caption Before attack                     Refer to caption Refer to caption Refer to caption Refer to caption After attack
@@
(a) MNIST Net1 Net2 Net3 Net4 Net5 Net1 1.000 0.644 0.646 0.642 0.576 Net2 X 1.000 0.698 0.699 0.581 Net3 X X 1.000 0.733 0.575 Net4 X X X 1.000 0.572 Net5 X X X X 1.000 (b) Fashion-MNIST Net1 Net2 Net3 Net4 Net5 Net1 1.000 0.200 0.207 0.214 0.167 Net2 X 1.000 0.278 0.301 0.208 Net3 X X 1.000 0.289 0.204 Net4 X X X 1.000 0.221 Net5 X X X X 1.000
(c) Kuzushiji-MNIST Net1 Net2 Net3 Net4 Net5 Net1 1.000 0.564 0.577 0.590 0.523 Net2 X 1.000 0.613 0.620 0.527 Net3 X X 1.000 0.671 0.544 Net4 X X X 1.000 0.510 Net5 X X X X 1.000 (d) Afro-MNIST Net1 Net2 Net3 Net4 Net5 Net1 1.000 0.621 0.613 0.628 0.465 Net2 X 1.000 0.700 0.712 0.551 Net3 X X 1.000 0.736 0.556 Net4 X X X 1.000 0.547 Net5 X X X X 1.000
@@
4.1 A. Small neural networks
We evaluated our method on five simple fully connected networks (FCNs) and a two-layer convolutional neural network (CNN). The number of neurons considered for each hidden layer is described in Table 1. These networks were trained on four datasets including MNIST mnist, Fashion-MNIST fashion, Kuzushiji-MNIST kuz, and Afro-MNIST afro. All these datasets are comprised of 60000 training examples and 10000 test examples. Each example consists of a $28\times 28$ image paired with a classification label from a set of 10 categories. MNIST, Fashion-MNIST, Kuzushiji-MNIST, and Afro-MNIST contain images of handwritten numerals, fashion, Hiragana characters, and African numbers respectively. The selection of datasets with similar attributes was deliberate as our focus was on understanding the interplay between networks, irrespective of data size or category.
@@
Table 4 depicts the robustness of FCNs against the black-box attack detailed in Algorithm 1. The impact of the attack on some test images is illustrated in Figure 1. Results for CNN are also provided, showcasing its inferior robustness, given its role in generating the adversaries. In our specific scenario, we selected a norm value of $\epsilon=\frac{10}{255}$ and $N_{ep}=10$. The results indicate that, on average, Net5, Net2, Net3, Net4, and Net1 ranked from vulnerable to robust. Among the FCNs, Net5 exhibited the lowest overall accuracy. Notably, according to Table 2, Net5 displayed the weakest correlation with the other networks, aligning with its subpar performance and larger number of layers. Conversely, Net4 demonstrated the highest average correlation score with the other networks, potentially attributed to its larger number of neurons per layer.
@@
4.2 B. Large neural networks
The proposed correlation metric was also investigated on large neural networks such as ResNet resnet, DenseNet densenet, and EfficientNet tan2019efficientnet. ResNet, short for Residual Network, is a deep neural network architecture that leverages residual blocks and skip connections to enhance training efficiency and performance. The variants of ResNet include ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 resnet. DenseNet, short for Dense Convolutional Network, is a deep neural network architecture that leverages dense connections between layers to promote feature reuse and efficient gradient flow, resulting in compact models that excel in various computer vision tasks. The architecture has several variants, namely DenseNet-121, DenseNet-161, DenseNet-169, and DenseNet-201 densenet. EfficientNet scales width, depth, and resolution proportionally to build efficient models (B0-B7) that achieve better accuracy with fewer parameters than other CNN models tan2019efficientnet. For all cases, we utilized the pretrained weights on ImageNet imagenet as our baselines. We limited our analysis to just 10 test data points sampled from the validation set of ImageNet due to the time and space complexities inherent in our methodology. Since investigating the correlation on all of the neurons in all layers was not possible, we focused our analysis solely on the initial output of the fourth layer for all ResNets, the primary output of the third transition layer for all DenseNets, and output of the third stage for all EfficientNets. A point to be highlighted is that assessing the correlation metric on the final layers often yields superior accuracy compared to other layers, given that these layers encapsulate more profound and meaningful representations. The results shown in Table 5 suggest that the most similar networks to ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 are ResNet-34, ResNet-18, ResNet-34, ResNet-34, and ResNet-34 respectively. As demonstrated in Table 6, DenseNet-161, DenseNet-121, DenseNet-201, and DenseNet-121 have the most similarities to DenseNet-121, DenseNet-161, DenseNet-169, and DenseNet-201. Similarly, Table 7 shows high correlation between adjacent scales. This pattern illustrates that networks with roughly the same number of layers often exhibit more profound partial correlations, serving as a validation of the effectiveness of our proposed correlation metric. Furthermore, deeper neural networks exhibit greater robustness against adversarial attacks. Thus, our findings further reveal that a stronger correlation with a robust model correlates with higher robustness.
@@
Table 7: Partial correlation between EfficientNets.
EfficientNet-B0 EfficientNet-B1 EfficientNet-B2 EfficientNet-B3 EfficientNet-B4
EfficientNet-B0 1 0.820 0.811 0.816 0.819
EfficientNet-B1 X 1 0.826 0.820 0.821
EfficientNet-B2 X X 1 0.819 0.814
EfficientNet-B3 X X X 1 0.822
EfficientNet-B4 X X X X 1
5 Discussion
@@
6 Conclusion
We introduced a simple neuronal correlation score that compares a candidate model to a well audited reference using only forward activations on a small, unlabeled probe set. The score matches units across models with a depth aware penalty and averages the result symmetrically. Across MNIST family networks and ImageNet pretrained architectures, higher representational alignment coincides with favorable validity indicators and with robustness trends under transferable attacks. These findings support the use of representational alignment as an external, training data independent signal that complements developer controlled evaluations.